OpenStudy (anonymous):

How can there be an x and y intercept of (0,0) if there is an asymptote at y=0?

OpenStudy (anonymous):

For example, in this rational function\[\frac{ x }{ x^2-1 }\] there is a horizontal asymptote at y=0, but there is still the intercept of (0,0). How can this be?

geerky42 (geerky42):

I suggest you to try and graph it

OpenStudy (phi):

It's not a straight line. It curves

OpenStudy (anonymous):

As you can see from the graph, it passes through the origin, yet there is an asymptote at y=0. I don't understand how this can happen. http://www.wolframalpha.com/input/?i=%28x%29%2F%28x%5E2-1%29

OpenStudy (phi):

you have \[ y = \frac{x}{x^2 - 1} \] when x is zero you get \[ y= \frac{0}{-1} = 0 \] when x is a very large number, x^2 - 1 is (almost) x^2 (compare 100^2 to 100^2 -1 , they are close in value) so for large positive x we can say x/(x^2-1) is about x/x^2 = 1/x and as you know for large x, 1/x approaches 0 i.e. y ->0 as x ->infinity so this function does go through (0,0) and it asymptotes to y=0 that is what non-linear expressions do: complicated things

OpenStudy (anonymous):

okay, thank you for that explanation!